# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v5

@@ 4 Feature modeling
[…]
with Dirac-delta function $\delta$, and mapping $F:\mathcal{X}\rightarrow\mathcal{Y}$. This relationship is characterized by the ambiguity in $\mathbf{X}$ and is inherently probabilistic, due to various reasons such as noise in the data (occlusions, sensor noise, insufficient resolution, etc.) or variability within a class (e.g. not all cats have tails). Hence, the observance of substantial aleatoric uncertainty can in some cases be inevitable, but may also signal the need for higher-quality data acquisition or shifting to another modality. The possible input dependency of the uncertainty develops into further categorization of either heteroscedastic (dependent) or homoscedastic (independent) aleatoric uncertainty. In most practical scenarios, aleatoric uncertainty modeling methods encompass both types and assume a parameterized likelihood function $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$ as a direct reflection of $p(\mathbf{Y}|\mathbf{X})$. For example, it is possible to model a distribution parameterized by the output of a CNN. Also, conditional generative models are used to learn the data distribution through so-called “latent” (i.e. unobserved) variables. For the purpose of taxonomy, we categorize models based on the location of these latent variables: those introduced near the output are discussed in Section 4.1, while lower-dimensional latents embedded deeper within the architecture are covered in Section 4.2. While their theoretical formulations are largely similar, their practical implications can differ significantly.
@@ 4.1 Pixel-level sampling
Uncertainty in segmentation masks can be modeled directly at the pixel level. These approaches can be further categorized into those that assume independence between pixels (Section 4.1.1), and those that explicitly model spatial correlations (Section 4.1.2). In the former case, accurate uncertainty estimates rely on well-calibrated models, which is an assumption that often fails in practice without additional tuning on a separate calibration set. In the latter case, a stochastic variable is introduced to capture dependencies between neighboring pixels.
@@ 4.1.1 Independence
Figure 4: Visualization of a reliability diagram, which illustrates whether the model’s predicted confidences align with the observed empirical accuracies.
As discussed earlier, neural network predictions are often normalized with SoftMax activation in order to interpret the confidence values as parameters of a probability mass distribution, but rerely reflect the true probabilities in modern neural networks Guo et al. (2017). Thus, interpreting the confidences as true probabilities is often only justified after proper validation, which is referred to as model calibration. Here, it is measured whether the empirical accuracy of a model approximately equals the provided class confidence $c_{k}$ for class $k$, i.e. $P(Y=k\,|\,c_{k})=c_{k}$. Calibration is typically visualized with a reliability diagram, where miscalibration and under-/over-confidence can be assessed by inspecting the deviation from the graph diagonal (Figure 4). Furthermore, different methods can be used to quantify calibration, although each may introduce its own biases. A fairly straightforward metric, the Expected Calibration Error (ECE), determines the normalized distance between accuracy (acc) and confidence (conf) bins as
with $n_{b}$ the number of samples in bin $b$ and $N$ being the total sample size across all bins. The ECE is prone to skew representations if some bins are significantly more populated with samples due to over-/under-confidence. Furthermore, the Maximum Calibration Error (MCE) is more appropriate for high-risk applications, where only the worst bin is considered.Additionally, when background pixels have a predominant influence, each bin can be weighted equally using the Average Calibration Error (ACE) to mitigate this imbalance Jungo et al. (2020); Neumann et al. (2018).
Nonetheless, contemporary neural networks often exhibit poorly calibrated uncertainty estimates. This misalignment is hypothesized due to the use of negative log-likelihood as the training objective, along with regularization techniques such as batch normalization, weight decay, and others (Guo et al., 2017). As a result, calibration methods are typically required to adjust the predicted confidences. Since most of these techniques are post-hoc, i.e. applied after training, they necessitate a separate validation set. For example, Temperature Scaling (Guo et al., 2017) has been applied in a pixel-wise manner for segmentation problems (Ding et al., 2021). Nonetheless, some methods, such as Label Smoothing (Silva & Oliveira, 2021; Liu et al., 2022a) or using the Focal Loss (Mukhoti et al., 2020) can be directly applied on the training data. Furthermore, over-fitting has often been considered to be the cause of over-confidence (Szegedy et al., 2016; Pereyra et al., 2017) and erroneous pixels can therefore be penalized through regularizing low-entropy outputs (Larrazabal et al., 2021a).
@@ 4.1.2 Spatial correlation
Figure 5: Illustration of the PixelCNN-based PixelSeg (Zhang et al., 2022c). Parameter ‘C’ indicates the concatenation module, ’R’ the resampling module and $\sigma$ the softmax activation. Dotted elements appear during test-time sampling.
Assuming pixel independence inhibits spatial coherence in uncertainty quantification, which is a key factor for segmentation. Introducing spatial correlation can be achieved through an autoregressive approach. For instance, we can rephrase Equation (7) to
@@
where pixel $Y_{i}$ is predicted based on the preceding pixels. A popular implementation of this formulation is known as the PixelCNN (Van Den Oord et al., 2016). For dense predictions, Zhang et al. (2022c) propose PixelSeg, which predicts a downsampled segmentation mask $\mathbf{\tilde{y}}$ with $p_{\bm{\phi}}(\mathbf{\tilde{y}}|\mathbf{x}$), and fuse this with a conventional CNN to predict the full-resolution mask with $p_{\bm{\theta}}(\mathbf{y}|\mathbf{\tilde{y}},\mathbf{x}$). The two masks are fused through a resampling module, containing a series of specific transformations to improve quality and diversity of the samples. Figure 5 illustrates this concept. Notably, PixelCNNs employ a recursive sampling process, which also enables completion/inpainting of user-given inputs.
Figure 6: Depiction of Stochastic Segmentation Networks (Monteiro et al., 2020). Here, the covariance of the likelihood distribution is explicitly modeled through a low-rank approximation.
@@ 4.2 Latent-level sampling
The limitation of modeling intricate, complex distributions directly at the output level can be mitigated by using generative models, which often rely on simpler, lower-dimensional latent variables $\mathbf{Z}\sim p_{\mathbf{Z}}$ with $\mathcal{Z}\in\mathbb{R}^{d}$, to instead learn the approximate through
@@
with parameters $\bm{\theta},\bm{\psi}$. As such, the spatial correlation is induced through mapping the latent variables to segmentation masks. Conditioning the latent density, i.e. $p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})$, on the input images is not a necessity, but usually preferred for smooth optimization trajectories Zheng et al. (2022). However, it is also possible to simply employ an unconditional prior $p(\mathbf{z})$. As argued by Kahl et al. (2024), their predictive uncertainty can be decomposed as
In contrast to Equation (3), the mutual information term here encapsulates the aleatoric uncertainty, while the expected entropy reflects the epistemic uncertainty. Also, note that the mutual information is between the output and latent variables, rather than the parameters of the model. The key overarching contribution of latent-level sampling methods lies in the conditioning of generative models on the input image. This section briefly introduces Generative Adversarial Networks (Section 4.2.1), Variational Autoencoders (Section 4.2.2), and Denoising Diffusion Probabilistic Models (Section 4.2.3), and outlines how these architectures are commonly adapted for probabilistic segmentation.
@@ 4.2.1 Generative Adversarial Networks (GANs)
[…]
@@
Figure 7: Diagram of the Calibrated Adversarial Refinement network Kassapis et al. (2021), based on the Generative Adversarial Network with additional loss terms.
While conditional GANs have been used for semantic segmentation earlier (Isola et al., 2017), Kassapis et al. (2021) explicitly contextualized the architecture within aleatoric uncertainty quantification, using their proposed Calibrated Adversarial Refinement (CAR) network (see Figure 7). The calibration network, $F_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$, initially provides a SoftMax activated prediction as $F_{\bm{\theta}}(\mathbf{x})=\*c$, with (cross-entropy) reconstruction loss
@@
[…]
@@
with hyperparameter $\lambda\geq 0$. The purpose of the calibration network is argued to be threefold. Namely, it (1) sets a calibration target for $\mathcal{L}_{\text{cal}}$, (2) provides an alternate representation of $\mathbf{X}$ to $G_{\bm{\phi}}$, and (3) allows for sample-free aleatoric uncertainty quantification. The refinement network can be seen as modeling the spatial dependency across the pixels, which enables sampling coherent segmentation maps.
@@ 4.2.2 Variational Autoencoders (VAEs)
$p(\mathbf{\mathbf{Y}|\mathbf{X}})$ $=\log\int q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})}\mathrm{d}\mathbf{z}$ (20a)
$\geq\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})}\right]$ (20b)
$=\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\bm{\phi}}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\operatorname{KL}[\,q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})||q_{\bm{\psi}}(\mathbf{z}|\mathbf{X})\,],$ (20c)
where Jensen’s inequality justifies moving the expectation outside the logarithm. The first term in Equation (20) represents the reconstruction cost of the decoder, subject to the latent code $\mathbf{Z}$ and input image $\mathbf{X}$. The second term is the Kullback-Leibler (KL) divergence between the approximate posterior and prior density. As a consequence of the mean-field approximation, all involved densities are modeled by axis-aligned Gaussian densities and amortized through neural networks, parameterized by $\bm{\phi}$, $\bm{\theta}$ and $\bm{\psi}$. The predictive distribution after observing dataset $\mathcal{D}$ is then obtained as
@@
The implementation of the conditional ELBO in Equation (20) can be achieved through the well known VAE architecture (Kingma & Welling, 2013). Furthermore, some additional design choices lead to the Probabilistic U-Net (PU-Net) (Kohl et al., 2019b). Firstly, the latent variable $\mathbf{Z}$ is only introduced at the final layers of a U-Net conditioned on the images. The latent vector is up-scaled through tiling and then concatenated with the feature maps of the penultimate layer, which is followed by a sequence of 1$\times$1 convolution for classification. When involving conditional latent variables in this manner, it is expected that most of the semantic feature extraction and delineation are performed in the U-Net, while the variability in the latent samples is almost exclusively related to the segmentation variability. Therefore, relatively smaller values of $d$ are feasible than what is commonly used in conventional image generation tasks. Similar to related research on the VAE (Zhao et al., 2017; Bousquet et al., 2017; Higgins et al., 2017; Van Den Oord et al., 2017; Rezende & Mohamed, 2015), much work has been dedicated to improving the PU-Net.
Figure 8: Depiction of the Probabilistic U-Net (Kohl et al., 2019b) based on a conditional Variational Autoencoder (Kingma & Welling, 2013). The latent samples are inserted at the final stages of a U-Net through a tiling operation.
@@
[…]
@@
where $\mathbf{z}_{k}$ and $\mathbf{z}_{k-1}$ are intermediate variables from intermediate densities and $\mathbf{z}_{0}=\mathbf{f}^{-1}(\mathbf{x})$. Equation (22) can be substituted in the conditional ELBO objective in Equation (20) to obtain
@@
where the objective consists of a reconstruction term, sample-based KL-divergence and a likelihood correction term for the change in probability density induced by the NF. Bhat et al. (2023; 2022a) compare this approach with other parameterizations of the latent space, including a mixture of Gaussians and low-rank approximation of the full covariance matrix. Valiuddin et al. (2024b) show that the latent space can converge to contain non-informative latent dimensions, undermining the capabilities of the latent-variable approach, generally referred to as mode or posterior collapse (Chen et al., 2016; Zhao et al., 2017). Their proposition considers the alternative formulation of the ELBO, specified as
$\log p(\mathbf{Y}|\mathbf{X})$ $\geq\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}\left[\log p_{\bm{\phi}}(\mathbf{Y}|\mathbf{X},\mathbf{z})\right]$ (24)
$\hskip 35.0pt-\operatorname{KL}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})\,||\,q_{\bm{\psi}}(\mathbf{z}|\mathbf{X})]-I(\mathbf{Y},\mathbf{Z}|\mathbf{X}).$
The proposed novel objective maximizes the contribution of the (expected) mutual information between latent and output variables, i.e. the aleatoric uncertainty term of the predictive entropy in Equation (13). This enables the introduction of the updated objective
$\mathcal{L}$ $=-\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\log p_{\phi}(\mathbf{Y}|\mathbf{X},\mathbf{z})]$ (25)
$\hskip 40.0pt+\alpha\cdot\operatorname{KL}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})]+\beta\cdot\operatorname{S}_{\epsilon}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})],$
with $\operatorname{S}_{\epsilon}$ being the Sinkhorn divergence Cuturi (2013) and $\alpha,\beta$ denoting hyperparameters. This adaptation results in a more uniform latent space leading to increased model performance. Also, modeling the ELBO of the joint density has been explored (Zhang et al., 2022b). This formulation results in an additional reconstruction term and forces the latent variables to be more congruent with the data. Furthermore, constraining the latent space to be discrete has resulted in some improvements, where it is hypothesized that this partially addresses the model collapse phenomenon (Qiu & Lui, 2020).
@@
$q_{\theta}(\mathbf{Z}_{1:T}|\mathbf{Y},\mathbf{X})=q_{\theta}(\mathbf{Z}_{1}|\mathbf{Y},\mathbf{X})\prod^{T}_{t=2}q_{\theta}(\mathbf{Z}_{t}|\mathbf{Z}_{t-1}).$ (26)
@@
$\mathcal{L}_{\text{GECO}}=\lambda$ $\cdot(\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\kappa)$ (28)
@@
Hyperparameter $\lambda$ is the Lagrange multiplier update through the Exponential Moving Average of the reconstruction, which is constrained to reach target value $\kappa$, empirically set beforehand to an appropriate value. Finally, online negative hard mining is used to only backpropagate 2% of the worst performing pixels, which are stochastically picked with the Gumbel-SoftMax trick (Jang et al., 2016; Huijben et al., 2022). Furthermore, PHiSeg (Baumgartner et al., 2019) takes a similar approach to the HPU-Net. However, PHiSeg places residual connections between latent vectors across decompositions rather than in the convolutional layers. Furthermore, deep supervision at each decomposition scale to enforce the disentanglement between latent variables.
Figure 9: Hierarchical Probabilistic U-Net (Kohl et al., 2019a) based on a hierarchical Variational Autoencoder (Kingma & Welling, 2013; Sønderby et al., 2016; Klushyn et al., 2019). Instead of a single latent code, multiple decomposition scales encode the segmentation variability, depending on the depth of the U-Net structure.

### v8

@@ 4 Feature modeling
Following Equation (2), we focus in this section on approaches that place uncertainty at the level of features. In most practical scenarios, methods assume a parameterized likelihood function $p(\mathbf{Y}|\mathbf{X},\bm{\theta})$, where parameters are point estimates. Accordingly, the central design decision is where to introduce auxiliary stochastic variables. Hence, we categorize models based on the location of these latent variables: those introduced near the output are discussed in Section 4.1, while lower-dimensional or latents embedded deeper within the architecture are covered in Section 4.2. Although their theoretical formulations are similar, the practical consequences can differ substantially.
@@ 4.1 Pixel-level sampling
Uncertainty in segmentation masks can be modeled directly at the pixel level. These approaches can be further categorized into those that assume independence between pixels (Section 4.1.1), and those that explicitly model spatial correlations (Section 4.1.2). In the former case, accurate uncertainty estimates rely on well-calibrated models, which is an assumption that often fails in practice without additional tuning on a separate calibration set. Furthermore, such methods do not address the challenge of spatial coherence. In contrast, the latter approach introduces a stochastic variable to capture dependencies between neighboring pixels.
@@ 4.1.1 Independence
While it is well established that assuming independence across pixels is problematic, it was a standard approach in much of the early literature. As discussed demonstrated in Equation (6), neural network predictions are often normalized with SoftMax activation in order to interpret the confidence values as parameters of a probability mass distribution. However, these values rarely reflect the true probabilities in modern neural networks, hypothesized due to the use of negative log-likelihood as the training objective, along with regularization techniques such as batch normalization, weight decay, and others (Guo et al., 2017). Thus, interpreting the confidences as true probabilities is often only justified after proper validation, which is referred to as model calibration.
Calibration
Different methods can be used to quantify whether the empirical accuracy of a model approximately equals the provided class confidence $c_{k}$ for class $k$, i.e. $P(Y=k\,|\,c_{k})=c_{k}$. A fairly straightforward metric, the Expected Calibration Error (ECE), determines the normalized distance between accuracy (acc) and confidence (conf) bins as
with $n_{b}$ the number of samples in bin $b$ and $N$ being the total sample size across all bins. The ECE is prone to skew representations if some bins are significantly more populated with samples due to over-/under-confidence. Furthermore, the Maximum Calibration Error (MCE) is more appropriate for high-risk applications, where only the worst bin is considered. Additionally, when background pixels have a predominant influence, each bin can be weighted equally using the Average Calibration Error (ACE) to mitigate this imbalance (Jungo et al., 2020; Neumann et al., 2018). Calibration is typically visualized with a reliability diagram, where miscalibration and under-/over-confidence can be assessed by inspecting the deviation from the graph diagonal (Figure 4).
@@
Figure 4: Visualization of model calibration with a reliability diagram, which demonstrates whether the predicted model confidences align with the observed empirical accuracies.
Methods
Most of the calibration techniques are post-hoc, i.e. applied after training, they necessitate a separate validation set. For example. over-fitting has often been considered to be the cause of over-confidence (Szegedy et al., 2016; Pereyra et al., 2017) and erroneous pixels can therefore be penalized through regularizing low-entropy outputs (Larrazabal et al., 2021b). Another prominent method is Temperature Scaling (Guo et al., 2017), which has been successfully applied in a pixel-wise manner to segmentation problems (Ding et al., 2021). The standard softmax probability in Equation (6) is modified as
$p(\mathbf{Y}=k\,|\,\mathbf{x}^{*},\bm{\theta})=\frac{e^{\mathbf{a}_{k}/T}}{\sum_{k}e^{\mathbf{a}_{k}/T}},$ (9)
with some scalar $T$ optimized on a validation set. For overconfident models, $T>1$ softens the estimated distribution, pushing output probabilities away from the extremes of 0 and 1 toward more moderate values. In contrast, label Smoothing (Silva and Oliveira, 2021; Liu et al., 2022a) require only modifying the training data. Instead of one-hot encoding $\mathbf{Y}$ in to $[0,1]$, a small positive random variable $\epsilon$ is introduced, which converts the targets into $\epsilon,1-\epsilon]$, discouraging the model from producing overly confident logit outputs. The Focal Loss (Mukhoti et al., 2020) adds a modulating factor to down-weight the loss from well-classified examples. This forces the model to focus more on hard-to-classify pixels, which can implicitly improve calibration by preventing it from becoming overly certain about easy examples.
Ultimately, relying on raw model outputs as probabilities typically yields sample-level spatial incoherence. Because modern neural networks are often uncalibrated by default, they require post-hoc calibration. Crucially, resolving spatial incoherence does not preclude good calibration, i.e., spatially correlated models can still be well calibrated. Conversely, calibration alone cannot impose spatial structure. Well-calibrated models may remain spatially incoherent. This asymmetry motivates placing greater emphasis on modeling spatial coherence rather than calibration alone.
@@ 4.1.2 Spatial correlation
Figure 5: Illustration of the PixelCNN-based PixelSeg (Zhang et al., 2022d). Parameter ‘C’ indicates the concatenation module, ’R’ the resampling module and $\sigma$ the softmax activation. Dotted elements appear during test-time sampling.
Introducing spatial correlation can be achieved through an autoregressive approach. For instance, we can rephrase Equation (7) to
@@
where pixel $Y_{i}$ is predicted based on the preceding pixels. A popular implementation of this formulation is known as the PixelCNN (Van Den Oord et al., 2016). For dense predictions, Zhang et al. (2022d) propose PixelSeg, which predicts a downsampled segmentation mask $\mathbf{\tilde{y}}$ with $p_{\bm{\phi}}(\mathbf{\tilde{y}}|\mathbf{x}$), and fuse this with a conventional CNN to predict the full-resolution mask with $p_{\bm{\theta}}(\mathbf{y}|\mathbf{\tilde{y}},\mathbf{x}$). The two masks are fused through a resampling module, containing a series of specific transformations to improve quality and diversity of the samples (Figure 5). Notably, the recursive sampling process also enables completion/inpainting of user-given inputs.
@@
Figure 6: Depiction of a Stochastic Segmentation Network (Monteiro et al., 2020). Here, the covariance of the likelihood distribution is explicitly modeled through a low-rank approximation.
Key takeaways: Calibrating standard segmentation models can produce valid confidence values. However, it still preserves the incorrect pixel-independence assumption and requires a separate validation set. A possible solution is to model the correlation directly in pixel space through autoregressive modeling or a low-rank approximation of the output distribution.
@@ 4.2 Latent-level sampling
The limitation of modeling intricate, complex distributions directly at the output level can be mitigated by using generative models, which often rely on simpler, lower-dimensional latent variables $\mathbf{Z}\sim p_{Z}$ with $\mathcal{Z}\in\mathbb{R}^{d}$, to instead learn the approximation through
@@
with parameters $\bm{\theta},\bm{\psi}$. As such, the spatial correlation is induced through mapping the latent variables to segmentation masks. On a related note, an unconditional prior $p(\mathbf{z})$ can also be employed. However, conditioning the latent density, i.e. $p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})$, is usually preferred for smooth optimization trajectories (Zheng et al., 2022). The key overarching contribution of latent-level sampling methods lies in the conditioning of generative models on the input image. This section briefly introduces Generative Adversarial Networks (Section 4.2.1), Variational Autoencoders (Section 4.2.2), and Denoising Diffusion Probabilistic Models (Section 4.2.3), and outlines how these architectures are commonly adapted for probabilistic segmentation.
@@ 4.2.1 Generative Adversarial Networks (GANs)
A straightforward approach is to simply learn the marginalization in Equation (12) through sampling from an unconditional prior density, $p_{Z}$, and mapping this to segmentation $\mathbf{Y}$ through a generator $G_{\bm{\phi}}:\mathcal{X}\times\mathcal{Z}\rightarrow\mathcal{Y}$.  Goodfellow et al. (2014) show that this approach can be notably enhanced through the incorporation of a discriminative function (the discriminator), denoted as $D_{\bm{\psi}}:\mathbb{R}^{C\times D}\rightarrow[0,1]$. In this way, $G_{\bm{\phi}}$ learns to reconstruct realistic images using the discriminative capabilities of $D_{\bm{\psi}}$, making sufficient guidance from $D_{\bm{\psi}}$ to $G_{\bm{\phi}}$ imperative. We can denote the cost of $G_{\bm{\phi}}$ in the GAN as the negative cost of $D_{\bm{\psi}}$ as
@@
Conditional GANs have been used for semantic segmentation in early literature (Isola et al., 2017). However, Kassapis et al. (2021) explicitly contextualized the architecture within uncertainty quantification, using their proposed Calibrated Adversarial Refinement (CAR) network (see Figure 7). The calibration network, $F_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$, initially provides a SoftMax activated prediction as $F_{\bm{\theta}}(\mathbf{x})=\*c$, with (cross-entropy) reconstruction loss
@@
$\mathcal{L}_{D}=$- $\mathbb{E}_{p_{\mathbf{Z}}}\mathbb{E}_{p_{\mathcal{D}}}[1-\log D_{\bm{\psi}}(G_{\bm{\phi}}(F_{\bm{\theta}}(\mathbf{x}),\mathbf{z}),\mathbf{x})\,]-\mathbb{E}_{p_{\mathcal{D}}}[\,\log D_{\bm{\psi}}(\mathbf{y})\,].$ (16)
@@
with hyperparameter $\lambda\geq 0$. The purpose of the calibration network is argued to be threefold. Namely, it (1) sets a calibration target for $\mathcal{L}_{\text{cal}}$, (2) provides an alternate representation of $\mathbf{X}$ to $G_{\bm{\phi}}$, and (3) allows for sample-free uncertainty quantification. The refinement network can be seen as modeling the spatial dependency across the pixels, which enables sampling coherent segmentation maps.
Figure 7: Diagram of the Calibrated Adversarial Refinement network (Kassapis et al., 2021), based on the Generative Adversarial Network (Goodfellow et al., 2014) with additional mechanisms and loss terms.
@@ 4.2.2 Variational Autoencoders (VAEs)
$\log p(\mathbf{\mathbf{Y}|\mathbf{X}})$ $=\log\int q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})}\mathrm{d}\mathbf{z}$
$\geq\int q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})\log\frac{p(\mathbf{z},\mathbf{Y}|\mathbf{X})}{q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})}\mathrm{d}\mathbf{z}$
$=\mathbb{E}_{q_{\bm{\theta}}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\bm{\phi}}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\operatorname{KL}[\,q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})||q_{\bm{\psi}}(\mathbf{z}|\mathbf{X})\,],$ (19)
where Jensen’s inequality justifies moving the logarithm inside the integral. The first term in Equation (19) represents the reconstruction cost of the decoder, subject to the latent code $\mathbf{Z}$ and input image $\mathbf{X}$. The second term is the Kullback-Leibler (KL) divergence between the approximate posterior and prior density. As a consequence of the mean-field approximation, all involved densities are modeled by axis-aligned gaussian densities and amortized through neural networks, parameterized by $\bm{\phi}$, $\bm{\theta}$ and $\bm{\psi}$. The predictive distribution after observing dataset $\mathcal{D}$ is then obtained as
@@
The implementation of the conditional ELBO in Equation (19) can be achieved through the well known VAE architecture (Kingma and Welling, 2013). Furthermore, some additional design choices lead to the segmentation-specific Probabilistic U-Net (PU-Net) (Kohl et al., 2019b). Firstly, the latent variable $\mathbf{Z}$ is only introduced at the final layers of a U-Net conditioned on the images. The latent vector is up-scaled through tiling and then concatenated with the feature maps of the penultimate layer, which is followed by a sequence of $\times1$ convolution for classification. When involving conditional latent variables in this manner, it is expected that most of the semantic feature extraction and delineation are performed in the U-Net, while the variability in the latent samples is almost exclusively related to the segmentation variability. Therefore, relatively smaller values of $d$ are feasible than what is commonly used in conventional image generation tasks. Similar to related research on the VAE (Zhao et al., 2017; Bousquet et al., 2017; Higgins et al., 2017; Van Den Oord et al., 2017; Rezende and Mohamed, 2015), much work has been dedicated to improving the PU-Net.
Figure 8: Depiction of the Probabilistic U-Net (Kohl et al., 2019b) based on a conditional Variational Autoencoder (Kingma and Welling, 2013). The latent samples are inserted at the final stages of a U-Net through a tiling operation.
@@
Augmenting a Normalizing Flow (NF) to the posterior density of a VAE is a commonly used tactic to improve its expressiveness (Rezende and Mohamed, 2015). This phenomenon has also been successfully applied to VAE-like models such as the PU-Net (Valiuddin et al., 2021; Selvan et al., 2020). NFs are a class of generative models that utilize $k$ consecutive bijective transformations $f_{k}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ as $f=f_{K}\circ\ldots\circ f_{k}\circ\ldots\circ f_{1}$, to express exact log-likelihoods of arbitrarily complex approximations of the posterior $p(\mathbf{z}|\mathbf{x})$, where symbol $\circ$ indicates functional composition. The approximated posteriors in turn induce an approximate of $p(\mathbf{x}|\mathbf{z})=p(\mathbf{z}|\mathbf{x})\frac{p(\mathbf{x})}{p(\mathbf{z})}$ and are often denoted as $\log p(\mathbf{x})$ for simplicity. The fractional term accounts for the change in probability space and corresponds to the log-Jacobian determinant, as given by the Change of Variables theorem. Thus, we can denote
@@
where $\mathbf{z}_{k}$ and $\mathbf{z}_{k-1}$ are variables from intermediate densities and $\mathbf{z}_{0}=\mathbf{f}^{-1}(\mathbf{x})$. Equation (21) can be substituted in the conditional ELBO objective in Equation (19) to obtain
@@
where the objective consists of a reconstruction term, sample-based KL-divergence and a likelihood correction term for the change in probability density induced by the NF. Bhat et al. (2023; 2022a) compare this approach with other parameterizations of the latent space, including a mixture of Gaussians and low-rank approximation of the full covariance matrix. Valiuddin et al. (2024b) show that the latent space can converge to contain non-informative latent dimensions, undermining the capabilities of the latent-variable approach, generally referred to as mode or posterior collapse (Chen et al., 2016; Zhao et al., 2017). Their proposition considers a modified objective
$\mathcal{L}$ $=-\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\log p_{\phi}(\mathbf{Y}|\mathbf{X},\mathbf{z})]+\alpha\cdot\operatorname{KL}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{Y},\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})]+(\beta-\alpha)\cdot\operatorname{S}_{\epsilon}[q_{\bm{\theta}}(\mathbf{z}|\mathbf{X})\,||\,p_{\bm{\psi}}(\mathbf{z}|\mathbf{X})],$ (23)
with $\operatorname{S}_{\epsilon}$ being the Sinkhorn divergence (Cuturi, 2013) and $\alpha,\beta$ denoting hyperparameters. This adaptation results in a more uniform latent space leading to increased model performance. Also, modeling the ELBO of the joint density has been explored (Zhang et al., 2022c). This formulation results in an additional reconstruction term and forces the latent variables to be more congruent with the data. Furthermore, constraining the latent space to be discrete has resulted in some improvements, where it is hypothesized that this partially addresses the model collapse phenomenon (Qiu and Lui, 2020).
@@
$q_{\theta}(\mathbf{Z}_{1:T}|\mathbf{Y},\mathbf{X})=q_{\theta}(\mathbf{Z}_{1}|\mathbf{Y},\mathbf{X})\prod^{T}_{t=2}q_{\theta}(\mathbf{Z}_{t}|\mathbf{Z}_{t-1},\mathbf{X}).$ (24)
@@
$\mathcal{L}_{\text{GECO}}=\lambda$ $\cdot||\mathbb{E}_{q_{\theta}(\mathbf{Z}|\mathbf{Y},\mathbf{X})}[\,\log p_{\phi}(\mathbf{Y}|\mathbf{z},\mathbf{X})\,]-\kappa||_{1}$ (26)
@@
Hyperparameter $\lambda$ is the Lagrange multiplier update through the Exponential Moving Average of the reconstruction, which is constrained to reach target value $\kappa$, empirically set beforehand to an appropriate value. Finally, online negative hard mining is used to only backpropagate 2% of the worst performing pixels, which are stochastically picked with the Gumbel-SoftMax trick (Jang et al., 2016; Huijben et al., 2022). Furthermore, PHiSeg (Baumgartner et al., 2019) takes a similar approach to the HPU-Net. However, PHiSeg places residual connections between latent vectors across decompositions rather than in the convolutional layers. Furthermore, deep supervision is placed at each decomposition scale to enforce disentanglement between the latent variables.
Figure 9: Hierarchical Probabilistic U-Net (Kohl et al., 2019a) based on a hierarchical Variational Autoencoder (Kingma and Welling, 2013; Sønderby et al., 2016; Klushyn et al., 2019). Instead of a single latent code, multiple decomposition scales encode the segmentation variability.